let f(x)=x+2 and g(x)= 1-x^2. Which of the following properly expresses g(f(x+1))?
a. -x^2+6x+10
b. -x^2-6x-8
c. -x^2+x+3
d. x^2-6x+8
e. x^2-x+10

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

let f(x)=x+2 and g(x)= 1-x^2. Which of the following properly expresses g(f(x+1))?
-----------
g[f(x+1)] = g[x+1+2} = g[x+3]
= 1-(x+3)^2
= 1-[x^2+6x+9]
= -x^2-6x-8
===============
Cheers,
Stan H.
===============
a. -x^2+6x+10
b. -x^2-6x-8
c. -x^2+x+3
d. x^2-6x+8
e. x^2-x+10

Answer by mananth(16946)

let f(x)=x+2 and g(x)= 1-x^2. Which of the following properly expresses g(f(x+1))?
option
f(x+1)=x+1+2
=x+3
...
g(x)=1-x^2
g(f(x+1)=1-(x+3)^2
1-(x^2+6x+9)
1-x^2-6x-9
-x^2-6x-8 option B
...
m.ananth@hotmail.ca